A is a 4.4 kg block and B is a 2.6 kg block. The coefficients of static and kinetic friction between A and the table are 0.18 and 0.15. 
a) determine the minimum mass of block C that must be placed on A to prevent it from sliding
b) block C is suddenly raised from A. What is the acceleration of block A?
Answer #1

2.6x9 = (4.4+ me legals 2.6 = 4.4 the )x0.18 to me =lo kogo Mog- MAMAG= (MBtmaja -0-15 *4.4x9.8= 7xa a=2.72 m/s2